OpenStudy (anonymous):

If f(x) = 2[to the power of x] + 2[to the power of negative x], find f(1), f(2) and f(-2). Is f(-6)= f(6)

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

f(1) = 3/2

OpenStudy (anonymous):

f(1)=2+1/2=5/2 f(2)=4+1/4=17/4 f(-2)=1/4+4=17/4

OpenStudy (anonymous):

the function is even so for all x f(x)=x(-x) so f(6)=f(-6)

OpenStudy (anonymous):

him is wrong here..

OpenStudy (anonymous):

How did you get that answer though? :L

OpenStudy (anonymous):

doing...

OpenStudy (anonymous):

\[f(x)=2^x+2^{-x}\]\[f(x)=2^x+\frac{1}{2^x}\]now put the values of x..

OpenStudy (anonymous):

So basically I just have to substitute the 'x' values for the equation, and that's it?

OpenStudy (anonymous):

\[f(-x)=2^{-x}+2^{-(-x)}=2^{-x}+2^x=f(x)\]

OpenStudy (anonymous):

yes...for example..

OpenStudy (anonymous):

\[f(1)=2^1+\frac{1}{2^1}=2+\frac12=\frac52\]

OpenStudy (anonymous):

Oh, I get it. That's all you need to do? ._.